OpenStudy (anonymous):

Will give medal! Solve -7/3n=-42 What is n?

OpenStudy (anonymous):

n=0.05

OpenStudy (anonymous):

It has to be a fraction:)

OpenStudy (anonymous):

1/18

OpenStudy (aum):

Is the n in the denominator or numerator? -7/3 * n or -7 / (3n) ?

OpenStudy (anonymous):

-7/-42 = 3n 1/6 x 1/3 = n n= 1/18

OpenStudy (anonymous):

Thank you! :)

OpenStudy (anonymous):

I am so sorry it needs to be a whole number

OpenStudy (aum):

If you answered my previous question I can help. Which one is the question? \[ -\frac{7}{3}n = -42 ~~~~~\text{OR} ~~~~ -\frac{7}{3n} = -42 \]

OpenStudy (anonymous):

the first one?

OpenStudy (aum):

-7/3 * n = -42 multiply by -1: 7/3 * n = 42 multiply by 3: 7n = 126 divide by 7: n = 126/7 = 18

OpenStudy (anonymous):

Can u help me with another one?

OpenStudy (aum):

I can try one more.

OpenStudy (anonymous):

Carla's plant was 10 inches tall. Yolanda's was 8 inches tall. How much taller was Carla's plant than Yolanda's? A. 1 in. B. 2 in. C. 2 in. D. 2

OpenStudy (aum):

Carla's plant was 10 inches tall. Yolanda's was 8 inches tall. How much taller was Carla's plant than Yolanda's? What are your thoughts?